OpenStudy (anonymous):

can any body show me how to express this in simplest form? x^2+9x+14/x^2-49 \[\div\] 3x+6/x^2+x-56

OpenStudy (anonymous):

we need to factor as much as possible

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

by factoring you mean?

OpenStudy (anonymous):

x^2+9x+14 factors to (x+2)(x+7)

OpenStudy (anonymous):

okay i get it

OpenStudy (anonymous):

x^2-49 factors to (x-7)(x+7)

OpenStudy (anonymous):

((x+2)(x+7))/((x-7)(x+7)) * (x-7)(x+8)/(3x+6)

OpenStudy (anonymous):

DONE!

OpenStudy (anonymous):

3x+6 factors to 3(x+2)

OpenStudy (anonymous):

x^2+x-56 factors to (x-7)(x+8)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

x^2+9x+14/x^2-49 = (x+2)/(x-7) 3x+6/x^2+x-56 = (3 (x+2))/((x-7) (x+8)) (x+2)/(x-7) * ((x-7) (x+8))/(3 (x+2)) (x+8)/3

OpenStudy (anonymous):

thats the simplified answer pk51?

OpenStudy (anonymous):

Yes, do you understand how I got it?

OpenStudy (anonymous):

so you factored both equations? then divided?

OpenStudy (anonymous):

Yes, but when dividing fractions, you flip one and the multiply across.

OpenStudy (anonymous):

no multiplied

OpenStudy (anonymous):

ohhhh!!!! okay i got it